| | For example, the probability of a tossed fair coin coming up heads is ½ = 0.5 = 50% because there is one favorable outcome (heads) out of 2 possible outcomes (heads and tails). Another common example is standard deck of 52 playing cards: there are 13 cards in each suit: spades, hearts, clubs and diamonds. After the deck is shuffled thoroughly, the probability that the top card on the deck is a heart is <math>13 / 52 = 1 / 4 = 0.20</math>, or 20%. | | For example, the probability of a tossed fair coin coming up heads is ½ = 0.5 = 50% because there is one favorable outcome (heads) out of 2 possible outcomes (heads and tails). Another common example is standard deck of 52 playing cards: there are 13 cards in each suit: spades, hearts, clubs and diamonds. After the deck is shuffled thoroughly, the probability that the top card on the deck is a heart is <math>13 / 52 = 1 / 4 = 0.20</math>, or 20%. |
| | + | Empirical (or experimental) probability is calculated simply as the ratio of ''observed'' favourable outcomes to total samples taken. Thus if a pair of dice might be rolled a thousand times, with an observed occurrence of 26 pairs of sizes. This gives an empirical probability of 26/1000 or 2.60%. With a larger number of samples, the empirical probability would be expected to tend towards the theoretical value. If it does not, then either the experimental [[methodology]] or the theoretical calculation and its assumptions would be questioned. In the case of the dice, the most likely explanation would be that the dice in the experiment were not truly unbiased.<ref>http://regentsprep.org/Regents/math/ALGEBRA/APR5/theoProp.htm</ref> |
| | + | In situations more complex than a pair of dice, the basis of theoretical probability calculations can become increasingly contentious, and the approach tends to be useful only in the case of limited or absent observational data. For example, when the outbreak of a new strain of [[influenza]] occurs, theoretical calculations of expected mortality rates are quickly produced based on extrapolation from other flu strains. These are soon superseded by actual mortality observations.<ref>http://seedmagazine.com/content/article/overestimating_avian_flu/</ref> |
| − | The famous mathematician Sir Frederick Hoyle once calculated that the chances of life occurring by chance are 1 in 10 to the power of 50, or 1 in 100,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000. | + | The famous mathematician Sir Frederick Hoyle once calculated that the chances of life occurring (i.e. the theoretical probability) by chance are 1 in 10 to the power of 50, or 1 in 100,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000. |